OpenStudy (anonymous):

What is the value of the discriminant? 4x2-7x+3=0

OpenStudy (turingtest):

for equations of the form\[ax^2+bx+c=0\]the discriminant is\[b^2-4ac\]

OpenStudy (anonymous):

What's the value of it?

OpenStudy (turingtest):

compare:\[ax^2+bx+c=0\]and you equation\[4x^2-7x+3=0\]what are a, b, abd c in your case? plug them into the formula. this site is not about giving direct answers. give it a try and show your work up until if/when you get stuck.

OpenStudy (turingtest):

your*

OpenStudy (turingtest):

and*

OpenStudy (anonymous):

you should be able to figure it out from the above posts. give it a try and show us what you get.